Calculate the grams of the indicated product when 17.9 g of the first reactant and 10.1 g of the second reactant is used:
Al2S3(s)+6H2O(l)→2Al(OH)3(aq)+3H2S(g) (H2S)
1
Expert's answer
2019-05-20T08:23:09-0400

Al2S3(s) + 6H2O(l) → 2Al(OH)3(aq) + 3H2S(g)


Calculating amounts of reactants:

"n(Al_2S_3) = \frac{m(Al_2S_3)}{M(Al_2S_3)} = \frac{17.9g}{(2\cdot 27 + 3 \cdot 32)\frac{g}{mol}}=\frac{17.9g}{150\frac{g}{mol}}=0.119mol"

"n(H_2O) = \frac{m(H_2O)}{M(H_2O)} = \frac{10.1g}{(2\cdot 1 + 16)\frac{g}{mol}}=\frac{10.1g}{18\frac{g}{mol}}=0.561mol"

According to stoichiometry of the reaction ratio Al2S3:H2O is 1:6. But according to given amounts of reagents this ratio is 0.119:0.561 = 1:4.7. This means that Al2S3 is excess reactant and H2O is limiting reactant.

Amount of Al2S3 reacted with water is:

"n(react. Al_2S_3)=\frac{1}{6}n(H_2O)=\frac{1}{6}\cdot 0.561mol = 0.0935 mol"

Amount of Al(OH)3 produced in the reaction:

"n(Al(OH)_3)=2n(react. Al_2S_3)=2\cdot 0.0935 mol = 0.187 mol"

"m(Al(OH)_3)=n(Al(OH)_3)\cdot M(Al(OH)_3) = 0.187mol \cdot 78 \frac{g}{mol}=14.6g"

Amount of H2S produced in the reaction:

"n(H_2S)=3n(react. Al_2S_3)=3\cdot 0.0935 mol = 0.2805mol"

"m(H_2S)=n(H_2S)\cdot M(H_2S) = 0.2805mol \cdot 34 \frac{g}{mol}=9.5g"